OpenStudy (anonymous):

Find a polynomial p of degree 3 such that -1, 3 and 6 are zeros of p and p(0)=1

OpenStudy (anonymous):

thought I had this one but my P(0) is 18. can you confirm that hthe problem is P(0) = 1

OpenStudy (anonymous):

I am positive that p(0)=1

OpenStudy (anonymous):

:(

Directrix (directrix):

y = 1/18(x+1)(x-3)(x-6)

Directrix (directrix):

y = (1/18) x^3 - (4/9) x^2 + (1/2) x + 1

OpenStudy (anonymous):

(x +1) (x -3) (x -6) (x +1) (x^2 - 9x + 18) x^3 - 9x^2 + 18x + (x^2 - 9x + 18) + 1 x^3 - 8x^2 + 9x + 19

OpenStudy (mani_jha):

Cant we do it, by taking a polynomial of form, \[ax ^{3}+bx ^{2}+cx+d\] and equating it to the various values given? I am getting rather odd values for a,b,c,d. But atleast the process is right

OpenStudy (dumbcow):

p(x) = a(x+1)(x-3)(x-6) p(0) = a(1)(-3)(-6) = 1 18a = 1 a = 1/18 p(x) = 1/18(x+1)(x-3)(x-6) = (1/18)x^3 - (4/9)x^2 +(1/2)x +1